Can someone please shoe me how to do this word problem with the steps?! I am stuck!
Given the simple interest formula A=P(1+Rt). , where A is the accumulated amount you will have in t years if you invest the principal P at the rate R.
a) Solve the formula for R. Leave the answer simplified.
b) What rate of interest are you earning if in 5 years your $6,000 investment has grown to be $7,140?

Answer by nyc_function(2741)

(a) Solve the formula for R. Leave the answer simplified.

We need to isolate R. In other words, we need to place R all alone on one side of the equation. Thisis what is meant by solve for R.
A = P(1 + Rt)
A = P + RPt
A - P = RPt
(A - P)/Pt = R
=========================================================
(b) What rate of interest are you earning if in 5 years your $6,000 investment has grown to be $7,140?
We need to use the given formula.
A = P(1 + Rt)
Let A = 7140
Let R = rate of interest
Let P = 6000
Let t = 5
We now plug into formula and simplify.
7140 = 6000(1 + R5)
7140 = 6000 + 30000R
7140 - 6000 = 30000R
1140 = 30000R
1140/30000 = R
0.038 = R
3.8% = R